# Adversarial Augmentation Training Makes Action Recognition Models More Robust to Realistic Video Distribution Shifts

### v1

@@ Abstract
Despite recent advances in video action recognition achieving strong performance on existing benchmarks, these models often lack robustness when faced with natural distribution shifts between training and test data. We propose two novel evaluation methods to assess model resilience to such distribution disparity. One method uses two different datasets collected from different sources and uses one for training and validation, and the other for testing. More precisely, we created dataset splits of HMDB-51 or UCF-101 for training, and Kinetics-400 for testing, using the subset of the classes that are overlapping in both train and test datasets. The other proposed method extracts the feature mean of each class from the target evaluation dataset’s training data (i.e. class prototype) and estimates test video prediction as a cosine similarity score between each sample to the class prototypes of each target class. This procedure does not alter model weights using the target dataset and it does not require aligning overlapping classes of two different datasets, thus is a very efficient method to test the model robustness to distribution shifts without prior knowledge of the target distribution. We address the robustness problem by adversarial augmentation training – generating augmented views of videos that are “hard” for the classification model by applying gradient ascent on the augmentation parameters – as well as “curriculum” scheduling the strength of the video augmentations. We experimentally demonstrate the superior performance of the proposed adversarial augmentation approach over baselines across three state-of-the-art action recognition models - TSM, Video Swin Transformer, and Uniformer. The presented work provides critical insight into model robustness to distribution shifts and presents effective techniques to enhance video action recognition performance in a real-world deployment.
@@ 1 Introduction
Data augmentation is one potential solution to account for this fragility. It is a popular method to create synthetic variations of the existing training data that will enable classification models to generalize better to previously unseen test data. However, it is not yet clear what kind of augmentation is necessary to generalize to test data with different distribution shifts. There has been much work on automatically selecting augmentation policies given training and validation data [8, 24, 35, 22, 34]. However, such approaches optimize augmentation of the training data, and it is not clear whether the resulting generalization applies to test data with much distribution shift.
@@
    We will publicly release the constructed subsets of HMDB-51, UCF-101, and Kinetics-400 to enable further research on this important problem.
@@ 2 Related Work
Data augmentation.    [45] summarizes image data augmentation techniques for deep learning. AutoAugment [8] is an augmentation policy search algorithm that finds the best augmentation on a target dataset, based on the highest validation accuracy. Due to AutoAugment’s expensive policy search, Population-Based Augmentation [24], FastAutoAugment [35], and FasterAutoAugment [22] proposed more efficient searching algorithms, by learning a schedule policy over a fixed-policy, using density matching, and using differential augmentation with a generative adversarial network (GAN) [18] architecture that involves a policy generator and a discriminator, respectively. Differentiable Automatic Data Augmentation [34] proposed a data augmentation policy searching algorithm using Relaxed Bernoulli distribution [26] which is differentiable, similar to FasterAutoaugment, and further introduced an unbiased gradient estimator that enables joint optimization of the augmentation policies and network parameters, instead of using a GAN. RandAugment [9] showed that simple random augmentations with randomly sampled transformations achieve similar performance more efficiently.
@@
These approaches were evaluated using simulated data, while we propose to use real data for testing. Evaluating robustness with augmented data prohibits the same augmentations to be used for training. This paper focuses on a more realistic scenario where a known set of data augmentation strategies is used for training and evaluating is done with unprocessed real data.
@@ 3 Problem and Methodology
To improve generalizability, the training data is augmented. Hard-to-classify adversarial examples are generated by applying gradient ascent on the augmentation parameters which are fully differentiable. We then train the classifier using AT loss, calculated using both clean and adversarial examples.
@@ 3.1 Adversarial Augmentation Training
$\begin{split}\mathcal{L_{\text{AT}}}(f_{\phi}(\mathbf{x}_{i}),f_{\phi}(\mathbf{x}^{\star}_{i}),y_{i})=&\alpha\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x}_{i}),y_{i})\\ &+(1-\alpha)\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x^{\star}}_{i}),y_{i})\end{split}$ (1)
@@
$\begin{split}\mathcal{L_{\text{AT-ME}}}(f_{\phi}(\mathbf{x}_{i}),f_{\phi}(\mathbf{x}^{\star}_{i}),y_{i})=&\alpha\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x}_{i}),y_{i})\\ &+(1-\alpha)\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x^{\star}}_{i}),y_{i})\\ &-\gamma\mathcal{L_{\text{E}}}(f_{\phi}(\mathbf{x^{\star}}_{i}))\end{split}$ (2)
@@ 3.2 Cross-Dataset Evaluation
Cosine Similarity Evaluation (Expt. 2).    The second method applies the feature extractor trained on the source dataset to the videos in the target dataset, which are split into a training subset and a test subset. This is a simpler method that does not require manual class matching. A more descriptive procedure can be found in Supplementary Material.
@@ 4 Experiments
HMDB/Kinetics and UCF/Kinetics Datasets (Expt. 1).    HMDB-51 [32] is a popular human action recognition dataset that is composed of around 7,000 video clips divided into 51 categories, collected from mainly movies as well as YouTube. The UCF-101 [49] dataset consists of 13,320 videos in 101 action classes collected from YouTube. Kinetics-400 [28] is a large-scale action dataset in which each video clip is around 10 seconds long, and there are over 300,000 videos in 400 classes.
@@
TruZe [21] is used to identify shared classes from the HMDB-51 and Kinetics-400, and UCF-101 and Kinetics-400 datasets, based on the visual and semantic similarity. More details on this procedure can be found in Supplementary Material. The final datasets are named HMDB-28/Kinetics-28 and UCF-65/Kinetics-65, where the training sets are subsets of the HMDB and UCF data, respectively, and test subsets come from Kinetics. The 28 and 65 refer to the number of shared classes. The three official published splits of HMDB-51 and UCF-101 are used, but only the shared classes are selected. The results in Table 1 are the average performance over the three splits.
HMDB normal-$\leftrightarrow$ UCF Evaluation (Expt. 2).    For the experiments using the cosine similarity function, the HMDB-28 trained models are tested on UCF-101, and the UCF-65 trained models are tested on HMDB-51, using the cosine similarity measure with class prototypes as predictions. The trained feature extractor is used to create prototype vectors from one part of the target dataset, and evaluation is based on the classification accuracy of the other part of the dataset. This assesses the quality of the feature extractor on another dataset with distribution shifts. The results in Table 1 are the average performance over the nine splits, three runs from the source dataset and each run evaluates with three splits from the target dataset.
@@
See Supplementary Material for implementation details.
@@ 5.1 Shared Class Experiments 1a, 1b
See Supplementary Material for confusion matrices that show per-class performance drop with distribution shifts and improvements using the proposed adversarial augmentation.
@@ 6 Conclusion
This paper addressed the problem of model generalization to realistic test distribution shifts. Two new datasets that are comprised of three existing datasets were created that shared the same subset of label classes. Although the same classes were used, the variety of videos in the original dataset sources meant that there was a huge distribution shift from the source to the target datasets. When using the target datasets, action recognition performance dropped significantly. This led to trying adversarial augmentation, with and without curriculum scheduling, as an approach to generating hard adversarially augmented videos. This approach gave a small but meaningful improvement in performance, even with the large distribution shift in the test data. The second cross-dataset evaluation approach, using the cosine similarity as logits, also showed a similar trend as the matching dataset experiments, providing a simpler alternative method without having to curate datasets with matching classes.

### v2

@@ Abstract
Despite recent advances in video action recognition achieving strong performance on existing benchmarks, these models often lack robustness when faced with natural distribution shifts between training and test data. We propose two novel evaluation methods to assess model resilience to such distribution disparity. One method uses two different datasets collected from different sources and uses one for training and validation, and the other for testing. More precisely, we created dataset splits of HMDB-51 or UCF-101 for training, and Kinetics-400 for testing, using the subset of the classes that are overlapping in both train and test datasets. The other proposed method extracts the feature mean of each class from the target evaluation dataset’s training data (i.e. class prototype), and estimates test video prediction as a cosine similarity score between each sample to the class prototypes of each target class. This procedure does not alter model weights using the target dataset and it does not require aligning overlapping classes of two different datasets, thus it is a very efficient method to test the model robustness to distribution shifts, without prior knowledge of the target distribution. We address the robustness problem by adversarial augmentation training – generating augmented views of videos that are “hard” for the classification model by applying gradient ascent on the augmentation parameters – as well as “curriculum” scheduling the strength of the video augmentations. We experimentally demonstrate the superior performance of the proposed adversarial augmentation approach over baselines across three state-of-the-art action recognition models - TSM, Video Swin Transformer, and Uniformer. The presented work provides critical insight into model robustness to distribution shifts and presents effective techniques to enhance video action recognition performance in a real-world deployment.
@@ 1 Introduction
Data augmentation is one potential solution to account for this fragility. It is a popular method to create synthetic variations of the existing training data that will enable classification models to generalize better to previously unseen test data. However, it is not yet clear what kind of augmentation is necessary to generalize to test data with different distribution shifts. There has been much work on automatically selecting augmentation policies given training and validation data [8, 24, 35, 22, 34]. However, such approaches optimize augmentation of the training data, and it is not clear how well the resulting generalization applies to test data with much distribution shift.
@@
    We publicly release the constructed subsets of HMDB-51, UCF-101, and Kinetics-400, along with all experimental code, to enable further research on this important problem. [^1] https://github.com/kiyoon/video-adversarial-augmentation
@@ 2 Related Work
Data augmentation.    [45] summarizes image data augmentation techniques for deep learning. AutoAugment [8] is an augmentation policy search algorithm that finds the best augmentation on a target dataset, based on the highest validation accuracy. Due to AutoAugment’s expensive policy search, Population-Based Augmentation [24], FastAutoAugment [35], and FasterAutoAugment [22] proposed more efficient searching algorithms, by learning a schedule policy over a fixed-policy, using density matching, and using differential augmentation with a generative adversarial network (GAN) [18] architecture that involves a policy generator and a discriminator, respectively. Differentiable Automatic Data Augmentation [34] proposed a data augmentation policy searching algorithm (using the Relaxed Bernoulli distribution [26]) which is differentiable, similar to FasterAutoaugment, and further introduced an unbiased gradient estimator that enables joint optimization of the augmentation policies and network parameters, instead of using a GAN. RandAugment [9] showed that simple random augmentations with randomly sampled transformations achieve similar performance more efficiently.
@@
These approaches were evaluated using simulated data, while we propose to use real data for testing. Evaluating robustness with augmented data prohibits the same augmentations to be used for training. This paper focuses on a more realistic scenario where a known set of data augmentation strategies is used for training and evaluation is done with unprocessed real data.
@@ 3 Problem and Methodology
To improve generalizability, the training data is augmented. Hard-to-classify adversarial examples are generated by applying gradient ascent on the augmentation parameters which are fully differentiable. We then train the classifier using the AT (Adversarial Training) loss, calculated using both clean and adversarial examples.
@@ 3.1 Adversarial Augmentation Training
$\mathcal{L_{\text{AT}}}(f_{\phi}(\mathbf{x}_{i}),f_{\phi}(\mathbf{x}^{\star}_{i}),y_{i})=\alpha\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x}_{i}),y_{i})+(1-\alpha)\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x^{\star}}_{i}),y_{i})$ (1)
@@
$\begin{split}\mathcal{L_{\text{AT-ME}}}(f_{\phi}(\mathbf{x}_{i}),f_{\phi}(\mathbf{x}^{\star}_{i}),y_{i})&=\alpha\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x}_{i}),y_{i})\\ &+(1-\alpha)\mathcal{L_{\text{CE}}}(f_{\phi}(\mathbf{x^{\star}}_{i}),y_{i})-\gamma\mathcal{L_{\text{E}}}(f_{\phi}(\mathbf{x^{\star}}_{i}))\end{split}$ (2)
@@ 3.2 Cross-Dataset Evaluation
Cosine Similarity Evaluation (Expt. 2).    The second method applies the feature extractor trained on the source dataset to the videos in the target dataset, which are split into a training subset and a test subset. This is a simpler method that does not require manual class matching. A more descriptive procedure can be found in the Supplementary Material.
@@ 4 Experiments
HMDB/Kinetics and UCF/Kinetics Datasets (Expt. 1).    HMDB-51 [32] is a popular human action recognition dataset that is composed of around 7,000 video clips divided into 51 categories, collected mainly from movies as well as YouTube. The UCF-101 [49] dataset consists of 13,320 videos in 101 action classes collected from YouTube. Kinetics-400 [28] is a large-scale action dataset in which each video clip is around 10 seconds long, and there are over 300,000 videos in 400 classes.
@@
TruZe [21] is used to identify shared classes from the HMDB-51 and Kinetics-400, and UCF-101 and Kinetics-400 datasets, based on the visual and semantic similarity. More details on this procedure can be found in the Supplementary Material. The final datasets are named HMDB-28/Kinetics-28 and UCF-65/Kinetics-65, where the training sets are subsets of the HMDB and UCF data, respectively, and test subsets come from Kinetics. The 28 and 65 refer to the number of shared classes. The three official published splits of HMDB-51 and UCF-101 are used, but only the shared classes are selected. The results in Table 1 are the average performance over the three splits. The HMDB-28/Kinetics-28 dataset consists of 3445 HMDB and 43406 Kinetics videos, and the UCF-65/Kinetics-65 dataset consists of 8935 UCF and 78583 Kinetics videos.
HMDB $\leftrightarrow$ UCF Evaluation (Expt. 2).    For the experiments using the cosine similarity function, the HMDB-28 trained models are tested on UCF-101, and the UCF-65 trained models are tested on HMDB-51, using the cosine similarity measure with class prototypes as predictions. The feature extractor trained using the training data is then used on the target dataset. It is used to create class prototype vectors (for each target class) from one part of the target dataset, and evaluation is based on the classification accuracy of the other part of the dataset. This assesses the quality of the feature extractor on another dataset with distribution shifts. The results in Table 1 are the average performance over the nine splits, three runs from the source dataset and each run evaluates with three splits from the target dataset.
@@
See the Supplementary Material for implementation details.
@@ 5.1 Shared Class Experiments 1a, 1b
See the Supplementary Material for confusion matrices that show per-class performance drop with distribution shifts and improvements using the proposed adversarial augmentation.
@@ 6 Conclusion
This paper addressed the problem of model generalization to realistic test distribution shifts. Two new datasets that are comprised of three existing datasets were created that share the same subset of label classes. Although the same classes were used, the variety of videos in the original dataset sources meant that there was a huge distribution shift from the source to the target datasets. When using the target datasets, action recognition performance dropped significantly. This led to trying adversarial augmentation, with and without curriculum scheduling, as an approach to generating hard adversarially augmented videos. This approach gave a small but meaningful improvement in performance, even with the large distribution shift in the test data. The second cross-dataset evaluation approach, using the cosine similarity as logits, also showed a similar trend as the matching dataset experiments, providing a simpler alternative method without having to curate datasets with matching classes.